In a certain store, there is a 0.05 probability that the scanned price in the bar code scanner will not match the advertised price. The cashier scans 850 items.

(a-1) What is the expected number of mismatches? (Round your answer to the nearest whole number.)
  Expected number   
(a-2)

What is the standard deviation? (Use your rounded number for the expected number of mismatches for the calculation of standard deviation. Round your final answer to 4 decimal places.)

  Standard deviation   
(b)

What is the probability of at least 37 mismatches? (Round the z-value to 2 decimal places. Use Appendix C-2 to find probabilities. Round your final answer to 4 decimal places.)

  Probability   
(c)

What is the probability of more than 49 mismatches? (Round the z-value to 2 decimal places. Use Appendix C-2 to find probabilities. Round your final answer to 4 decimal places.)

  Probability   
Answer #1

a-1)

expected number of mismatches =np =42.5

a-2)  Standard deviation =(np(1-p))1/2 =6.3541

b)

probability = P(X>36.5) = P(Z>-0.94)= 1-P(Z<-0.94)= 1-0.1736= 0.8264

c)

probability = P(X>49.5) = P(Z>1.1)= 1-P(Z<1.1)= 1-0.8643= 0.1357